OpenStudy (anonymous):

3x+2y=1 whats the slope ?

OpenStudy (anonymous):

-3/2

OpenStudy (anonymous):

subtract x from both sides and then divide by two. this makes m equal to -3/2

OpenStudy (nurali):

3x +2y = 1. 2y = -3x +1. y = 1/2(-3x + 1). y = -3/2x + 1/2. so slope is -3/2.

OpenStudy (jhannybean):

put it in slope intercept form y = mx+b, your "m" is the slope. subtract -3x from both sides 2y = -3x+1 divide both sides by 2 2y/2 = -3x/2 +1/2 y = -3/2 x+1/2